umerlodhi (umerlodhi):

Question 1.1. Simplify: 5(8 + 7) (Points : 1) 75 65 47 15 Question 2.2. Use the distributive property to rewrite the expression: 5 • 37 (Points : 1) 5 • 30 + 7 5 • 30 + 5 • 37 15 • 30 + 15 • 7 5 • 30 + 5 • 7 Question 3.3. Simplify: 6y + 5z + 8y – 3z (Points : 1) 11y + 5z 14y + 2z 14y + 8z 16yz Question 4.4. Which pair of terms is a pair of like terms? (Points : 1) g2h and gh2 3g and 3h g4 and h4 4g2 and 9g2 Question 5.5. Combine

umerlodhi (umerlodhi):

Combine the like terms to simplify the expression: 5x3 – 2x2 + x + 4x3 (Points : 1) 8x3 9x6– 2x2+ x 9x3– 2x2+ x 5x3– 2x2+ x + 4x3 (the expression cannot be simplified)

jagr2713 (jagr2713):

1. all you have to do is distribute and multiply 5*8=40 and 5*7=35 and add them =75 igtg sorry but i gaved you at least one of them

OpenStudy (anonymous):

1- A 2- A 3- B 4-D 5- C

OpenStudy (anonymous):

Question 1 : 5 (8 + 7) Step 1 : 8+7 = 15 Step 2 : 15 x 5 = 75 Your answer is 75 Question 2 : 5 x 37 Widen it out 5 x 30 + 7 Question 3 : 6y + 5z +8y - 3z Step 1 : 6y + 8y = 14 Step 2 : 5z - 3z = 2 Your answer = 14y + 2z Question 4 : Your answer is C

umerlodhi (umerlodhi):

thanks everybody